Detectives have discovered the murder victim in juja at 6:30 am and the body temperature of the victim was then 260celcius after 30minutes the police surgeon arrived and found the body temperature to be 230 Celsius. If the air temperature was 160 Celsius throughout and the normal body temperature is 370 Celsius at what time did the police surgeon estimate that the crime had occurred(victim murdered)?


1
Expert's answer
2021-08-04T13:11:05-0400

"\text{At 6:30 a.m, t = 0 and the temperature $T_0=26^0C$}
\\\text{After 30 minutes t = 0.5hr and the temperature $T_0=23^0C$ and the room temperature}\\\text{The change in temperature is given by }\\\frac{dT}{dt}=-k(T-T_c)=-k(T-16)\\\text{Integrating, we have that}\\In(T-16)=-kt+c\\\text{Taking exponents, we have}\\T-16=Ae^{-kt}\text{ where, $A=e^c$, therefore}\\T=Ae^{-kt}+16
\\\text{when t=0 and $T_0=26^0C$, we have that A = 10}\\\implies T=Ae^{-kt}+16
\\\text{when t=0.5 and $T_0=23^0C$, we have that k = 0.7133}\\\implies T=10e^{-0.7133t}+16
\\\text{To calculate the time of death, we take T as normal body temperature and we }\\\text{compute t}\\37=10e^{-0.7133t}+16
\\\text{Hence, $t=-1.040hr$, therefore the death occured about 1 hour before 6:30 a.m}"